OpenStudy (anonymous):

2,272 ounces= Pounds 142pounds 284pounds 322pounds none of the above

OpenStudy (anonymous):

1p=16 ounces

OpenStudy (anonymous):

\[\frac{1p}{?p}=\frac{16 oun}{2.272 oun}\] cross and time 2.272*1p =16 p=16/2.272